OpenStudy (idealist10):

Solve y'=(y^3+2xy^2+x^2*y+x^3)/(x(y+x)^2) implicitly.

OpenStudy (idealist10):

y=ux u=y/x y'=u+u'x u'x+u=((ux)^3+2x(ux)^2+x^2(ux)+x^3))/(x(ux+x)^2) u'x+u=((ux)^2(ux+2x)+x^2(ux+x))/(x(ux+x)^2) How do I simplify this? The answer in the book is (y+x)^3=3x^3(ln abs(x)+C). How do I get to the answer from there?

OpenStudy (idealist10):

@amistre64

OpenStudy (rational):

x cancels out everytime you substitute `y=ux` in a homogeneous equation

OpenStudy (rational):

u'x+u = ((ux)^3+2x(ux)^2+x^2(ux)+x^3)) / (x(ux+x)^2) first cancel x

OpenStudy (idealist10):

What do I get if I cancel x?

OpenStudy (rational):

you will get a separable equation

OpenStudy (rational):

thats the only idea behind y=ux substitution

OpenStudy (idealist10):

Let me factor out first. Please don't leave.

OpenStudy (rational):

okie

OpenStudy (idealist10):

I don't see how I can cancel out the x. From u'x+u=((ux)^3+2x(ux)^2+x^2(ux)+x^3)/(x(ux+x)^2) u'x+u=((ux)^2(ux+2x)+x^2(ux+x))/(x(ux+x)^2) u'x+u=(x(ux)^2(u+2)+x^2(ux+x))/(x(ux+x)^2)

OpenStudy (rational):

\[\large \begin{align}\\ u'x+u = \dfrac{(ux)^3+2x(ux)^2+x^2*(ux)+x^3}{x(ux+x)^2} \end{align}\]

OpenStudy (rational):

Notice that each term in the numerator of right hand side has a factor x^3

OpenStudy (rational):

\[\large \begin{align}\\ u'x+u &= \dfrac{(ux)^3+2x(ux)^2+x^2*(ux)+x^3}{x(ux+x)^2} \\~\\ &= \dfrac{x^3[u^3+2u^2+u+1]}{x^3(u+1)^2} \\~\\ \end{align}\]

OpenStudy (idealist10):

But how do I simplify (u^3+2u^2+u+1)/(u+1)^2 after I cancel out the x^3?

OpenStudy (rational):

not sure yet, but thats just an algebra problem

OpenStudy (rational):

maybe send that u to right side first

OpenStudy (rational):

\[\large \begin{align}\\ u'x+u &=\dfrac{u^3+2u^2+u+1}{(u+1)^2} \\~\\ u'x &=\dfrac{u^3+2u^2+u+1}{(u+1)^2} - u \\~\\ \end{align}\]

OpenStudy (rational):

get common denominator and simplify

OpenStudy (idealist10):

Got it! Thanks a lot!